to find A when P =$11,000.00 time is 5 year, rate is5.4% n is 12: and or 5.5% where n is 4

Answer by rfer(16322)

A=11000(1+0.054/12)^12*5
A=11000(1.0045)^60
A=11000*1.30917
A=$14,400.87
------------------
A=11000(1+0.055/4)^4*5
A=11000(1.01375)^20
A=11000*1.314
A=$14,454.00